Fill in the blanks with >,
45 – (– 11) ______ 57 + (– 4) 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

45–(-11) 45 + 11 = 56
57 + (-4) = 57–4 = 53
56 > 53
Therefore 45–(-11) > 57 + (-4)

Thanks